OpenStudy (vane11):

Can someone see if they get the same answer as me please? Use the laws of logarithms to expand and simplify the expression. ln(x^3/(sqrt(x) (1+x)^3)) (4/2)lnx-3ln(1+x)

OpenStudy (mathmale):

I'd be glad to give you feedback if only you'll post / share your own work here.

OpenStudy (zzr0ck3r):

\(ln(\frac{x^3}{\sqrt{x}(1+x^3)})=ln(x^3)-ln(\sqrt{x}*(1+x)^3)=ln(x^3)-ln(x^{\frac{1}{2}})-ln(1+x)^3=\\3ln(x)-\frac{1}{2}ln(x)-3ln(1+x)=\frac{5}{3}ln(x)-3ln(1+x)\)

OpenStudy (zzr0ck3r):

not sure how you got 4

OpenStudy (vane11):

oh wow I added the whole number 3 to the 1/2 hence the 4/2, I knew something was funky

OpenStudy (vane11):

Thank you!

OpenStudy (zzr0ck3r):

np